OpenStudy (anonymous):

Consider the functions f(x) = 2(x − 7) and g(x) = 2x. What is the value of (f + g)(4)? −6 14 2 −2

OpenStudy (anikhalder):

(f+g)(x)=f(x)+g(x) Hence, (f+g)(4)=f(4)+g(4) Now, f(4)=2(4-7) =>f(4)=-6 g(4)=2x4 =>g(4)=8 So, (f+g)(4)=f(4)+g(4) =>(f+g)(4)=-6+8 =>(f+g)(4)=2 Practise more here: http://www.purplemath.com/modules/fcnops.htm